OpenStudy (anonymous):

i don't get this one.. Use Guess, Check, and Revise to find values for x, y, and z that make all three equations true: x + 1 = y y - 4 = z 2z = x

OpenStudy (anonymous):

z = 1

OpenStudy (anonymous):

x = 1/2

OpenStudy (anonymous):

y = 5

OpenStudy (anonymous):

x = 2 (corrected)

OpenStudy (anonymous):

how did u get that?

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

ive been trying to figure this problem out for like 30 minutes

OpenStudy (anonymous):

z = 3 ; x = 6 ; y = 7

OpenStudy (anonymous):

put x = 2z and y = z+4 into x+1 = y

OpenStudy (anonymous):

then it will be 2z + 1 = z + 4

OpenStudy (anonymous):

So, z = 3

OpenStudy (anonymous):

You can do the rest

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yeah, thanx a LOT XD

OpenStudy (anonymous):

Good luck